OpenStudy (anonymous):

(3)/(4+4sqrt(5)) whenever I simplify this I end up with something different than the answer. I was hoping someone could explain the process to me quickly! Thanks

OpenStudy (anonymous):

\[3/(4+4\sqrt{5}) * ((4-4\sqrt{5}) /(4-4\sqrt{5}) = (12-12\sqrt{5})/(16-16*5) =3(1-\sqrt{5})/16\]

OpenStudy (anonymous):

Thanks so much I didn't even think about using the conjugate!